I have been trying and trying here and running in circles! I even have my brother on it: Any help would be greatly appreciated!!!!
- The cost of plywood sheeting recently increased by $1.39. per sheet. A contractor purchased several for a current project and spent $166.80 total. Six months ago, for a different project, he spent the same total but bought six more sheets than his current project. How many sheets of plywood did the contractor but for his current project?
I know that he spent a total of $8.34 more. I have x+6p=$166.80; "p" for plywood. I started dividing that into the total and ran into price of six.
This one is kicking me and I really don't think it should!! I am not even nearly on the right track! Any help would be awesome!! Thank you tons!! rl

Answer by CharlesG2(828)

I have been trying and trying here and running in circles! I even have my brother on it: Any help would be greatly appreciated!!!!
- The cost of plywood sheeting recently increased by $1.39. per sheet. A contractor purchased several for a current project and spent $166.80 total. Six months ago, for a different project, he spent the same total but bought six more sheets than his current project. How many sheets of plywood did the contractor but for his current project?
I know that he spent a total of $8.34 more. I have x+6p=$166.80; "p" for plywood. I started dividing that into the total and ran into price of six.
This one is kicking me and I really don't think it should!! I am not even nearly on the right track! Any help would be awesome!! Thank you tons!! rl
(p + 1.39)*x = 166.80
p * (x+6) = 166.80 (same total but x+6 or 6 more sheets)
(p + 1.39)*x = p * (x+6)
px + 1.39x = px + 6p
1.39x = 6p
x = 6p/1.39
(1.39/6)x = p
((1.39/6)x + 1.39)*x = 166.80
(1.39x^2)/6 + 1.39x = 166.80
(x^2)/6 + x = 120
x^2 + 6x = 720
x^2 + 6x - 720 = 0
(x+30)(x-24) = 0 (30*-24=-720)
x=24 since that is the positive answer
(1.39/6)x = p
(1.39/6)*24 = p
1.39 * 4 = p
5.56 = p
check:
(p + 1.39)*x = 166.80
(5.56 + 1.39) * 24 = 166.80
6.95 * 24 = 166.80
166.80 = 166.80
check:
p * (x+6) = 166.80 (24 + 6 = 30)
5.56 * 30 = 166.80
166.80 = 166.80
note:
am not sure where you got the $8.34